Find the magnitude of gravitational force a 61.8 kg person would experience while standing on the surface of Earth with a mass of 5.98*10^ 24 kg and a radius of 6.37 * 10 ^ 6 * m . The universal gravitational constant is 6.673*10^ -11 N* m^ 2 /kg^ 2 Answer in units of m



1
Expert's answer
2021-03-29T09:02:04-0400

The magnitude of gravitational force will be


"F=G\frac{Mm}{R^2},\\\space\\
F=6.673·10^{-11}\frac{5.98·10^{24}·61.8}{(6.37·10 ^ 6)^2}=608\text{ N}."